OpenStudy (anonymous):

Line L has equation 2x - 3y = 5. Line M passes through the point (2, -10) and is perpendicular to line L. Determine the equation for line M. A) 2x + 3y = -14 B) 2x - 3y = -26 C) 3x + 2y = -14 D) 3x - 2y = -26

OpenStudy (anonymous):

I undestand how to get the slope and everything, but I don't understand how to use the ordered pair it provides...

OpenStudy (ash2326):

@Sunshine447 what's the slope of line m?

OpenStudy (anonymous):

-3/2?

OpenStudy (ash2326):

Good, do you know how to find slope?

OpenStudy (anonymous):

using the slope formula and solving?

OpenStudy (ash2326):

yes

OpenStudy (anonymous):

okay, but how do I involve the ordered pair it gives me?

OpenStudy (ash2326):

Let line passes through (x, y) and (x1, y1) and has slope m, then equation is given as \[\frac{y-y1}{x-x1}=m\] here (x1, y1)=(2, -10) and slope is \(\frac{-3}{2}\)

OpenStudy (ash2326):

@Sunshine447 I think you can do now, can't you?

OpenStudy (anonymous):

i know its c based on the slope, but I don't understand where you get the -14 from

OpenStudy (anonymous):

@ash2326 u there?

OpenStudy (anonymous):

@Hero can you help me?

hero (hero):

Hint: The product of the slopes of two perpendicular lines is - 1

OpenStudy (anonymous):

I don't understand... How do you get the -14 from?

OpenStudy (anonymous):

*where do you

OpenStudy (anonymous):

ok put line l into the standard form.

OpenStudy (anonymous):

it is in standard form... isnt it?

OpenStudy (anonymous):

sorry i meant y=mx+b then when yo get it into that form you have to plug the points into the equation \[y-y _{1}=m \left( x-x _{1} \right)\]

OpenStudy (anonymous):

now the m is the slope from the first line and it has to be the reciprocal. so find that of the equation and plug it in.

OpenStudy (anonymous):

y=2/3x+5

OpenStudy (anonymous):

thats the mx+b right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok now turn the m into the reciprical

OpenStudy (anonymous):

opposite reciprocal or reciprocal?

OpenStudy (anonymous):

negative reciprocal

OpenStudy (anonymous):

-3/2

OpenStudy (anonymous):

ok goo now plug it into the m in to the formula i gave you

OpenStudy (anonymous):

is this point slope form?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ohhhhhh okay I have my formula sheet but I didn't know which one to use. It's c right?

OpenStudy (anonymous):

yeah i thinks so.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

haha a return of favors huh ?